# Ambiguity Clustering: an accurate and efficient decoder for qLDPC codes

### v1

@@
21 June 2024)
Error correction allows a quantum computer to preserve a state long beyond the decoherence time of its physical qubits by encoding logical qubits in a larger number of physical qubits. The leading proposal for a scheme of quantum error correction is based on the surface code, but several recently proposed quantum low-density parity check (qLDPC) codes allow more logical information to be encoded in significantly fewer physical qubits. Key to any scheme of quantum error correction is the decoder, an algorithm that estimates the error state of the qubits from the results of syndrome measurements performed on them. The surface code has a variety of fast and accurate decoders, but the state-of-the-art decoder for general qLDPC codes, BP-OSD, has a high computational complexity. Here we introduce Ambiguity Clustering (AC), an algorithm which seeks to divide the measurement data into clusters which are decoded independently. We benchmark AC on the recently proposed bivariate bicycle codes and find that, at physically realistic error rates, AC is between one and three orders of magnitude faster than BP-OSD with no reduction in logical fidelity. Our CPU implementation of AC is already fast enough to decode the 144-qubit Gross code in real time for neutral atom and trapped ion systems.
[…]
Quantum computing has the potential to enable breakthrough calculations in high impact fields including pharmaceutical [1] and material [2] simulation and cryptography [3]. Practical advantage in these areas, however, will only come with fault-tolerant quantum computing, where the constituent qubits of the computer are encoded with sufficient redundancy to endure the rapid and inevitable errors to which they are subject. Such a fault-tolerant system requires an encoding scheme—a quantum code—for the qubits, in which the state of some number of logical qubits is encoded in a larger number of physical qubits. The scheme also prescribes sets of measurements to be performed on the physical qubits, from which one can estimate whether an error has occurred and, if it has, determine an appropriate correction. This calculation is referred to as decoding. Here we describe Ambiguity Clustering, a new decoding algorithm applicable to arbitrary quantum low-density parity check (qLDPC) codes.
A quantum code is qLDPC if each prescribed measurement involves only a small number of physical qubits, and each physical qubit is involved in only a small number of measurements. These are natural conditions for any practical scheme of fault-tolerant computation. The most widely studied quantum code is Kitaev’s surface code [4]. In addition to being qLDPC, the pattern of measurement allows surface codes to be implemented on a flat surface using a square grid of qubits and nearest-neighbour interactions. This is a useful property for systems such as superconducting circuits where qubits are implemented as immovable circuit elements. The surface code also has excellent decoders. It was remarked in Kitaev’s original paper that decoding can be reduced to the problem of finding minimum-weight perfect matchings in a graph. This can be solved exactly in polynomial time by Edmonds’ Blossom algorithm [5], and more recently approximated in almost-linear time by the Union-Find decoder [6].
The surface code has a high overhead: a great number of physical qubits are required to encode a single logical qubit. Recently much work has focused on the development of more general qLDPC codes [7, 8, 9, 10]. These have the potential to encode many more logical qubits in a given set of physical qubits, whilst maintaining a similar level of protection. In return, we must [11] perform more complicated sets of measurements (not limited to local groups of physical qubits laid out on a flat surface) and we lose access to the efficient matching-based decoders.
The state of the art decoder for qLDPC codes is Belief Propagation followed by Ordered Statistics Decoding (BP-OSD) [12]. Belief Propagation (BP) is a classical inference algorithm which effectively decodes classical LDPC codes such as those used in the 5G network [13]. OSD is a post-processing step which handles the fact that quantum codes generally have multiple good explanations for any given set of measurement results (quantum degeneracy). BP-OSD suffers from a high computational complexity, as it involves both an expensive linear algebra step, cubic in the size of the system, and a search step which is in the worst case exponential in the size of the system.
In this paper we propose a new decoding algorithm, Ambiguity Clustering (AC), for decoding general qLDPC codes. AC works by forming clusters in a graph describing the code, guided by the output of BP. These clusters can be solved independently, so that the cost of the search step scales only with the size of the clusters, not with the overall size of the decoding problem. We compare BP-OSD and AC on Bravyi et al.’s recently proposed bivariate bicycle codes [14]. On these codes, simulated subject to a full circuit-level noise model, AC matches the accuracy of BP-OSD with one to three orders of magnitude of speedup, using Roffe’s standard implementation of BP-OSD [15, 16] as a benchmark.
Figure 1: Top. Logical error rates for quantum memory using the bivariate bicycle codes, labelled by code parameters $[[n,k,d]]$ (the number of physical qubits, logical qubits, and the distance of the code). Syndromes were extracted for $d$ rounds; logical error and timing data are per round of syndrome extraction. Dashed lines are the BP-OSD-CS(7) (see Section 3) data from [14]. For each data point, the AC parameter $K$ (Section 4.2) has been set to match or exceed the accuracy of BP-OSD. The shaded region indicates one (binomial) standard deviation of error. Only experiments for which we observed at least 5 failures are shown. The [[108,8,10]] code is omitted for clarity as it has similar performance to the [[90,8,10]] code.  Bottom. Decoding time per round of syndrome extraction at $p=0.3\%$. Bravyi et al. only obtained BP-OSD data down to $p=0.035$ for the $[[288,12,18]]$ code so for that code we set $K$ to match accuracy based on an extrapolation.
@@ 21 June 2024)
Figure 2: a. A Tanner graph for the 5-bit repetition code, with 5 error nodes (circles) and 4 check nodes (squares). Checks are each connected to an adjacent pair of error nodes, representing the condition that adjacent bits of the codewords $00000$ and $11111$ agree. An unknown error pattern (marked circles) has an observable syndrome (marked squares) consisting of the checks adjacent to an odd number of errors. Below is the associated parity check matrix $H$. The syndrome $\sigma$ is the mod 2 sum of the columns corresponding to the unknown error. Any solution $e$ to the equation $He=\sigma$ is an explanation of the observed syndrome.   b, c. A different choice of Tanner graph and the corresponding parity check matrix. The checks here represent the conditions that each bit of a codeword should agree with the final bit—an alternative way to ensure that all of the bits agree. This structure makes it particularly easy to read off the sets of columns summing to the observed syndrome: columns 2 and 3 (b), or column 5 together with columns 1 and 4 (c), corresponding to $e=(0,1,1,0,0)$ and $e=(1,0,0,1,1)$.  d. An arbitrary parity check matrix can be converted to the same special form by Gaussian elimination. Typically the parity check matrix is much wider than it is tall, so there are far more combinations of columns to search through to find the best solutions. BP-OSD streamlines this search by reordering the columns according to the output of BP, placing columns considered more likely to represent an error further to the left, before performing Gaussian elimination. High accuracy can then be obtained by searching over only small sets of columns from the right-hand side of the matrix.  e. Suppose that the parity check matrix has a partial block structure, covering the rows where the syndrome is non-zero, and that we’re confident that the rightmost block of columns corresponds to errors which did not occur. Then we can search over the possible explanations of each section of the syndrome separately. The part of the overall solution space explored by this search can be exponentially larger, in the number of blocks, than a search which does the same amount of work without regard to the block structure.  f. The blocks correspond to isolated clusters in an updated Tanner graph. The key idea of AC is that, by careful manipulation of the parity check matrix, this block structure can usually be obtained.
@@
Figure 2 gives an overview of the decoding problem and the main innovation of AC. Information about the errors inside a quantum computer is available only indirectly, in the form of a syndrome which indicates which of some set of parity checks are violated. Each syndrome can be explained by multiple different underlying errors. The decoding problem is to map observed syndromes to the most likely underlying errors.
@@
A decoding problem can be specified by either a Tanner graph or a parity check matrix. A Tanner graph has two types of nodes, check nodes and error nodes, with an edge between a check node and an error node whenever the corresponding check is sensitive to the corresponding error (Figure 2). The parity check matrix $H$ is the adjacency matrix of the Tanner graph: $H_{ij}$ is $1$ if the $i$th check is sensitive to the $j$th error, and 00 otherwise. Given a syndrome $\sigma$, we seek the most likely underlying errors $e$ such that
@@
The state of the art for this problem is BP-OSD.
BP is a message-passing algorithm which uses the observed syndrome to compute an estimate of how likely each individual error is to have occurred. We can turn this into a guess for the global error pattern $e$ by setting $e_{j}=1$ precisely when BP considers that the $j$th error is more likely to have occurred than not. This is, however, a poor decoder for quantum codes, typically not even producing solutions to (1).
In contrast, when the parity check matrix is in reduced form (Figures 22) it is easy to read off solutions to (1), but the number of solutions is typically exponentially large in the problem size. Ordered Statistics Decoding (OSD) uses the output of BP to direct the search to the most promising parts of the solution space.
@@
We first sort the columns of the parity check matrix according to the output of BP, so that columns corresponding to the most likely errors are on the left of the matrix. Then we perform Gaussian elimination (see Section 3.1), proceeding left to right. This converts the matrix to reduced form (Figure 2). If the output of BP is reliable, then the best solutions to (1) will largely comprise columns in the resulting identity block, so that accurate results can be obtained by considering only relatively few combinations of the remaining columns.
Our new decoder, AC, works by producing a block structure in the parity check matrix $H$ (Figure 2), or equivalently, a cluster structure in the Tanner graph (Figure 2). This structure is obtained by modifying the Gaussian elimination algorithm to choose its next step dynamically based on the syndrome as well as the output of BP, together with a new stopping condition that prevents it from running to completion. The advantage of this block/cluster structure is twofold. First, it is much cheaper to obtain than the fully reduced form in Figure 2. Second, we can use it to search over local solutions to each cluster independently.
The remainder of the paper is set out as follows. Section 1 introduces the decoding problem. Sections 2 and 3 set the stage for a description of AC by discussing Belief Propagation, Gaussian elimination, and how they are combined in BP-OSD. In Section 4, we describe AC in full. Section 5 details the Monte Carlo simulations used to produce Figure 1.
@@ 1 The decoding problem
The Ambiguity Clustering decoder can be applied to any problem expressed in terms of parity checks on probabilistically independent error mechanisms.
@@
The input to a decoding problem comprises
@@
The maximum likelihood decoder computes the logical effect vector $\lambda$ maximising the value of
@@
For a decoding problem to be well-posed, there must be at least one $e$ such that $He=\sigma$. We will assume that the rows of $H$ are linearly independent, which guarantees that we can always find such an $e$. This simplifies the presentation, but avoids no essential difficulties, provided $\sigma$ is in the column span of $H$.
Definition 1.1 models many different types of decoding problem, such as classical linear binary codes and quantum stabiliser codes with stochastic noise and perfect measurement. Here we focus instead on quantum memory with a circuit-level noise model. We are given a circuit which initialises a block of logical qubits, implements syndrome extraction for some number of rounds, then measures off the physical qubits in such a way that the value of the logical $Z$, say, operators can be determined. In this model, $H$ and $L$ have the following interpretations.
  • The errors represented by the columns of $H$ and $L$ are $n$ independent error mechanisms in the circuit, such as a particular single-qubit Pauli error on an idle qubit, a particular two-qubit Pauli error after a two-qubit gate, or an error in preparation or measurement. This $n$ will generally be many times larger than the number of physical qubits in the code.
  • The checks represented by the rows of $H$ are $m$ detectors: sums (mod 2) of physical measurement results which are deterministically 00 if no errors occur, such as measurements of the same stabiliser in consecutive rounds. The $ij$th entry of $H$ is $1$ if the value of the $i$th detector is flipped whenever the $j$th error occurs, and 00 otherwise.
  • The rows of $L$ correspond to $k$ logical observables whose values we are trying to preserve. The $ij$th entry of $L$ is $1$ if the value of the $i$th logical observable is flipped whenever the $j$th error occurs, and 00 otherwise.
For quantum error correction $n$ is typically much larger than $m+k$, meaning that, for any $e\in\mathbb{F}_{2}^{n}$, there are a very large number (at least $2^{n-m-k}$) of $e^{\prime}$ with both $He^{\prime}=He$ and $Le^{\prime}=Le$; that is, that have an identical syndrome and logical effect. This degeneracy is a key difference between quantum and classical LDPC codes.
2 Belief Propagation and the
split belief problem
Given an observed syndrome $\sigma$, what is the updated probability that $e_{j}=1$? Belief Propagation (BP) is an efficient algorithm to produce estimates $p_{j}$ for these posterior probabilities, which works by passing messages over the associated Tanner graph.
Classical decoding problems are non-degenerate. For each logical effect $\lambda$, there is a unique error $e$ with that logical effect and the observed syndrome. Provided the most likely solution is sufficiently likely, we can find it by rounding off the posterior probabilities:
This is not the case for quantum decoding problems. Even when there is an unambiguously most likely logical effect $\lambda$, that logical effect can be caused by many different physical errors with the observed syndrome. In this case, looking at the most likely value of each $e_{i}$ individually might not even produce an error $e$ compatible with the observed syndrome (see Figure 3). This is known as the split belief problem, and configurations leading to this situation as trapping sets [19, 20].
Figure 3: The split belief problem. The syndrome marked in red has four explanations a, b, c, d. If all six errors (circular nodes) have the same prior probability, then their posterior probabilities are identical (and close to $1/3$) (e). It is impossible to retrieve a solution to (1) by applying a uniform threshold to the posteriors.
Our new decoder, AC, builds upon the ideas of BP-OSD, which we describe in some detail in the next section.

### v2

@@
3 January 2025)
@@
Quantum computing has the potential to enable breakthrough calculations in high impact fields including pharmaceutical [1] and material [2] simulation and cryptography [3]. Practical advantage in these areas, however, will only come with fault-tolerant quantum computing, where the constituent qubits of the computer are encoded with sufficient redundancy to endure the rapid and inevitable errors to which they are subject. Such a fault-tolerant system requires an encoding scheme—a quantum code—for the qubits, in which the state of some number of logical qubits is encoded in a larger number of physical qubits. The scheme also prescribes sets of measurements to be performed on the physical qubits, from which one can estimate whether an error has occurred and, if it has, determine an appropriate correction, a calculation referred to as decoding. For practical reasons, it is preferred that each set of measurements involves only a small number of qubits, and each qubit is involved in only a small number of measurements. A quantum low-density parity check (qLDPC) code family is one for which these quantities are bounded by a constant as the number of physical qubits grows large.
[…]
Bravyi et al. [9] recently proposed implementing a code in the two-block group algebra (2BGA) family [10, 11] which can protect 12 logical qubits against $0.1\%$ circuit-level noise for one million rounds of syndrome extraction using 288 physical qubits. The surface code by contrast requires 241 physical qubits to provide only a single logical qubit with a slightly lower level of protection, a greater than 10×10$\times$10 × increase in the number of physical qubits per logical qubit. Avoiding the high qubit overhead of the surface code has motivated extensive research into more general qLDPC codes [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
There remain several obstacles to the widespread adoption of non-surface qLDPC codes. Proposals for computing on the encoded data [27, 28] are less well-developed than the lattice surgery protocols for the surface code [29]. The circuits implementing the codes will generally require more than nearest neighbour interactions on a square grid [30, 31]. We also lose access to efficient matching-based decoders.
Specialised decoders exist for certain families of codes [32, 33, 34, 35, 36], but the state of the art decoder for general qLDPC codes is Belief Propagation followed by Ordered Statistics Decoding (BP-OSD) [37]. Belief Propagation (BP) is a classical inference algorithm which effectively decodes classical LDPC codes such as those used in the 5G network [38]. OSD is a post-processing step which handles the fact that quantum codes generally have multiple good explanations for any given set of measurement results (quantum degeneracy). BP-OSD suffers from a high computational complexity, as it involves both an expensive linear algebra step, cubic in the size of the system, and a search step which is in the worst case exponential in the size of the system. Measurement data must be processed by the decoder as fast as it is generated, so a decoder which can’t keep up requires [39] elaborate ‘windowing’ schemes to process data non-sequentially and in parallel [40, 41], complicating overall system design. Additionally, slow decoders limit our ability to investigate new qLDPC codes.
In this paper we propose a new decoding algorithm, Ambiguity Clustering (AC), for decoding general qLDPC codes. AC works by forming clusters in a graph describing the code, guided by the (potentially ambiguous) output of BP. These clusters can be explored independently, so that the cost of the search step scales only with the size of the clusters, not with the overall size of the decoding problem.
Ambiguity Clustering
[…]
@@
Figure 1: The decoding problem and main innovation of AC. a. A Tanner graph for the 5-bit repetition code. Check nodes (squares) are each connected to an adjacent pair of error nodes (circles), representing the condition that adjacent bits of the codewords $00000$ and $11111$ agree. An unknown error pattern (shaded circles) has an observable syndrome (shaded squares) consisting of the checks adjacent to an odd number of errors. Below is the associated parity check matrix $H$. The syndrome $\sigma$ is the mod 2 sum of the columns corresponding to the unknown error. Any $e$ satisfying $He=\sigma$ is an explanation of the observed syndrome.   b, c. A different choice of Tanner graph and parity check matrix. The checks represent the conditions that each bit of a codeword should agree with the final bit. It is particularly easy to read off the sets of columns summing to the observed syndrome: columns 2 and 3 (b), or column 5 together with columns 1 and 4 (c), corresponding to $e=(0,1,1,0,0)$ and $e=(1,0,0,1,1)$.  d. An arbitrary parity check matrix (left) can be converted to a similar special form (right) by Gaussian elimination. There are typically a great many combinations of columns to search through. BP-OSD streamlines this search by reordering the columns according to the output of BP, placing columns more likely to represent an error further to the left, before performing Gaussian elimination. High accuracy can then be obtained by searching over only small sets of columns from the right-hand side of the matrix.  e. A partial block structure, covering the rows where the syndrome is non-zero, which would be useful for decoding. If we are confident that errors corresponding to the rightmost block of columns did not occur, then we can search over explanations for each section of the syndrome separately. The space explored by this search can be exponentially larger, in the number of blocks, than a search which does the same amount of work without regard to the block structure.  f. Blocks correspond to isolated clusters in an updated Tanner graph. The key idea of AC is that, by careful manipulation of the parity check matrix, this block structure can usually be obtained.
@@
Information about the errors inside a quantum computer is available only indirectly, in the form of a syndrome which indicates which of some set of parity checks are violated. Each syndrome can be explained by multiple different underlying errors. The decoding problem is that of mapping observed syndromes to the most likely underlying errors.
@@
Figure 1 gives an overview of the decoding problem and the main innovation of AC. A decoding problem can be specified by either a Tanner graph or a parity check matrix. A Tanner graph has two types of nodes, check nodes and error nodes, with an edge between a check node and an error node whenever the corresponding check is sensitive to the corresponding error (Figure 1). The parity check matrix $H$ is the adjacency matrix of the Tanner graph: $H_{ij}$ is $1$ if the $i$th check is sensitive to the $j$th error, and 00 otherwise. Given a syndrome $\sigma$, we seek the most likely underlying error vector $e$ such that
@@
Belief Propagation (BP) is an efficient algorithm to produce estimates $p_{j}$ for the updated posterior probability that $e_{j}=1$ (the $j$th error occurred), given (1). BP works by passing messages over the associated Tanner graph.
Classical decoding problems are non-degenerate, meaning there is typically only one likely solution to (1). Provided this solution is sufficiently likely, we can find it by rounding off the posterior probabilities:
This is not the case for quantum decoding problems. There may be many likely solutions to (1) with an identical effect on the stored logical information, so there is no need to distinguish between them. The degeneracy of quantum codes nonetheless poses a problem for BP-based decoders, as looking at the most likely value of each $e_{j}$ individually might not produce an error $e$ compatible with the observed syndrome (see Figure 2). This is known as the split belief problem, and configurations leading to this situation are known as trapping sets [42, 43].
Figure 2: The split belief problem. The syndrome marked in red has two explanations a, b. If all four errors (circular nodes) have the same prior probability, then their posterior probabilities are identical (in fact, equal to $1/2$) (c). It is impossible to retrieve a solution to (1) by applying a uniform threshold to the posteriors.
@@
Gaussian elimination (see Section 2.1) is a standard algorithm that converts a parity check matrix to reduced form (Figures 11). When a parity check matrix is in reduced form it is easy to read off solutions to (1), but the number of solutions is typically exponentially large in the problem size. Most of these solutions will be highly unlikely, requiring many more errors to occur during circuit execution than we expect. Without further guidance, finding likely solutions $e$ therefore quickly becomes prohibitively expensive as the problem size increases. BP with Ordered Statistics Decoding post-processing (BP-OSD) uses the output of BP to direct the search to the most promising parts of the solution space. In BP-OSD, we first sort the columns of the parity check matrix according to the output of BP, so that columns corresponding to the most likely errors are on the left of the matrix. Then we perform Gaussian elimination, proceeding left to right. This converts the matrix to reduced form. If the output of BP is reliable, then the best solutions to (1) will largely comprise columns in the resulting identity block, so that accurate results can be obtained by considering only relatively few combinations of the remaining columns.
Broadly speaking, AC works by performing incomplete Gaussian elimination on $H$, with modified pivot selection rules and modified stopping criteria. This has two advantages: first, it is cheaper than full Gaussian elimination. Second, the incomplete Gaussian elimination imposes a block structure on $H$, which allows for a more efficient search over the solution space. We emphasise that, as the parity check matrix $H$ is transformed, we can reinterpret it as a transformed Tanner graph, and the intuition for many of the steps of AC comes from thinking about this transformed Tanner graph.
After running BP to obtain estimated posteriors $p_{j}$, AC proceeds in three stages, which we detail in Section 3. In the first stage, we transform the Tanner graph to obtain an initial solution to $He=\sigma$. At every step of this first stage, we form a simple cluster consisting of just one check node and one error node. By the end of the stage, every marked check (corresponding to a $1$ in the syndrome) is in such a cluster. This allows us to find an initial solution by choosing, for each marked check, the error in its associated cluster.
In the second stage, we grow outwards from this initial solution to identify additional errors, not present in the initial solution, that could be relevant to solving the decoding problem. During this stage, we grow and merge the simple clusters formed in stage 1 into larger clusters. This gives us access to a larger space of solutions to look through for each cluster.
In the third stage, we analyse each of the clusters separately to find the likely effect the error in each cluster had on the stored logical information. We show that for most clusters this logical effect can be calculated very cheaply. For the remaining ambiguous clusters, we perform an approximate maximum-likelihood search to determine the most likely logical effect. Combining the logical effect of all the clusters gives us an estimate of the overall effect on the stored logical information.
Performance comparison
Figure 3: Comparison of BP-OSD and AC.Top. Logical error rates for quantum memory using the bivariate bicycle codes, labelled by code parameters $[[n,k,d]]$ (the number of physical qubits, logical qubits, and the distance of the code). Syndromes were extracted for $d$ rounds; logical error and timing data are per round of syndrome extraction. Dashed lines are the BP-OSD-CS(7) (see Section 2) data from [9]. For each data point, the AC parameter $K$ (Section 3.2) has been set to match or exceed the accuracy of BP-OSD. The shaded region indicates one (binomial) standard deviation of error. Only experiments for which we observed at least 5 failures are shown. The [[108,8,10]] code is omitted for clarity as it has similar performance to the [[90,8,10]] code.  Bottom. Decoding time per round of syndrome extraction at $p=0.3\%$ ($p=0.35\%$ for the [[288,12,18]] code).
We compare BP-OSD and AC on Bravyi et al.’s recently proposed bivariate bicycle codes [9]. These are 2BGA codes specified by pairs of two-variable polynomials. In addition to their low qubit overhead these codes were designed with a view to implementation on a superconducting quantum computer, as the required qubit connections could be implemented by making use of both sides of a 2D plane. Bravyi et al. present data for five codes with data qubit counts ranging from 72 to 288. When we configure AC to match the accuracy of BP-OSD on these codes, simulated subject to a full circuit-level noise model, we observe up to a 27×27$\times$27 × speedup, using Roffe’s standard implementation of BP-OSD [49, 50] as a benchmark.
[…]
Ambiguity Clustering represents a significant step towards real-time decoding of qLDPC codes on near future hardware, especially on slower qubit types including neutral atom systems. Real-time decoding on solid-state quantum systems, such as superconducting- and spin-based processors, requires a further 2–3 orders of magnitude improvement in decoding time over the CPU decoder presented here. Bridging this gap will likely require a combination of algorithmic advances and implementation on specialised computing hardware.
Related work
After this work was announced, Hillman et al. [53] announced the Localised Statistics Decoder (LSD), which also seeks to avoid using full Gaussian elimination to solve the decoding problem. AC is driven by global selection of the error most likely to help explain the syndrome at each step, with error clusters condensing from the set of considered errors as the algorithm proceeds. By contrast, LSD begins with a cluster of errors around each $1$ in the syndrome and is driven by local choices of the most likely adjacent error to pull into each cluster at each step. Rather than the single representation of the decoding matrix used by AC, LSD maintains independent copies of the parts of the matrix corresponding to each cluster, with rules for reconciling these representations as clusters grow into each other. We might say that, where AC is top-down, LSD is bottom-up, which could affect the available strategies for parallelising each algorithm. The existing, CPU, implementations of both AC and LSD are fully serial.
Acknowledgements
We would like to thank Mark Turner for numerous helpful discussions about software, and Dan Browne, Earl Campbell, Joan Camps, Ophelia Crawford, Maria Maragkou and Luigi Martiradonna for their valuable comments on drafts of this manuscript.
@@ 1 The decoding problem
Ambiguity Clustering can be applied to any problem expressed in terms of parity checks on probabilistically independent error mechanisms.
@@
A decoding problem comprises
@@
The maximum likelihood decoder computes the logical effect vector $\lambda$ maximising
@@
For a decoding problem to be well-posed, there must be at least one $e$ such that $He=\sigma$. To simplify the presentation, we will make the stronger assumption that the rows of $H$ are linearly independent, which guarantees that we can always find such an $e$. This assumption is not necessary (and the decoder implementations do not require it) but in any case can always be achieved by removing linearly dependent rows from $H$.
Definition 1.1 models many different types of decoding problem, such as classical linear binary codes and quantum stabiliser codes with stochastic noise and perfect measurement. Here we focus instead on quantum memory with a circuit-level noise model. See Section A.3 for a detailed description of the construction of $H$ and $L$ in this setting.
For quantum error correction $n$ is typically much larger than $m+k$, so for any $e\in\mathbb{F}_{2}^{n}$, there are a large number (at least $2^{n-m-k}$) of $e^{\prime}$ with both $He^{\prime}=He$ and $Le^{\prime}=Le$; that is, that have identical syndromes and logical effects. This degeneracy is a key difference between quantum and classical codes.